Solve for m and nwhere m and n are positive integer:2^m-2^n=48

Grade:12th Pass

1 Answers

Arun
6 years ago

totally based on power set concept

P(set1) = 2m

P(set2) = 2n

acc to given cond.. difference is 48.

2m - 2= 48

2m - 2n = 64 - 16

2m - 2n = 26 - 24

this implies 2= 26 2n = 24

this implies m = 6 n = 4